# High-frequency gravitational wave transients from superradiance

### 03 Apr 2026 v2

@@ 1 Introduction
where $\Omega_{H}$ is the angular velocity of the Kerr horizon and $m$ the azimuthal quantum number, is exponentially amplified at the expense of the black hole’s rotational energy [34, 35, 25]. When the boson’s Compton wavelength is comparable to the gravitational radius, the amplified modes are gravitationally trapped and grow into a macroscopic quasi-bound cloud—a gravitational atom (GA)—with occupation numbers reaching $N\sim 10^{70}$$0^{80} [11$, 9]. The dynamics are controlled by the gravitational fine-structure constant
@@
The GW frequency is set almost entirely by $\mu$: annihilation of two cloud bosons into a graviton emits at $\omega_{\rm ann}\simeq 2\mu$, while level transitions between states with principal quantum numbers $n_{g}$ and $n_{e}$ emit at the Bohr frequency $\omega_{\rm tr}\propto\mu\alpha^{2}(n_{g}^{-2}-n_{e}^{-2})$. For stellar-mass BHs this places signals in the audio band probed by LIGO/Virgo/KAGRA, and an extensive literature has explored these GW signatures and the spin constraints they imply [9, 4, 19, 22, Tsukada:2018]. A qualitatively distinct regime arises for primordial black holes (PBHs) [10, 26, 3]. For sub-solar mass PBH masses the superradiance condition selects boson masses $\mu\sim 10^{-7}$$10^{-5}$ eV, driving GW emission into the MHz–GHz band. This is precisely the frequency range targeted by resonant-cavity haloscopes such as ADMX, which were originally designed to search for axion dark matter but are equally sensitive to oscillatory strain at GHz frequencies, provided the signal persists for longer than the cavity ring-up time $\tau_{\rm ring}\sim 1/\Delta f_{\rm band}$. Also, notably, bosons in that mass range are independently motivated by the axion solution to the strong CP problem [23, 30, 31] and the cosmology thereof [24, 1, 13, 12].
@@
These signals have been studied in detail for isolated stellar-mass GAs, and we extend that framework here to the PBH mass range and the associated GHz frequency band [9, 4, 19]. Bosenova collapses, in which the cloud implodes once its occupation number exceeds a critical value set by bosonic self-interactions, introduce additional burst-like GW emission and periodically reset the superradiant growth [32, 33]; we account for the resulting upper bound on cloud populations when deriving strain amplitudes.
In realistic astrophysical environments, however, GAs are often not isolated. A binary companion introduces a time-varying tidal potential that drives resonant transitions between GA levels whenever the orbital frequency sweeps through a Bohr frequency of the gravitational atom. In this regime the cloud behaves as a driven two-level system, and its response is captured by a Landau–Zener analysis [7, 6]: depending on the adiabaticity parameter $z\equiv\eta^{2}/(|\Delta m|\dot{\Omega})$, where $\eta$ is the tidal mixing amplitude, transitions range from partial population transfer ($z\ll 1$) to complete level inversion ($z\gg 1$). The resulting GW signal is a short-duration transient burst whose characteristic strain and spectral shape have recently been computed by Kyriazis & Yang [20], whose formalism we adopt and apply to the PBH binary case. Binary-induced resonant transitions also imply:
@@ 2.2 Superradiant Instability Rates and Growth Timescales
$\Gamma_{n\ell m}\;\simeq\;\frac{C_{n\ell m}(a_{*})}{r_{g}}\;\alpha^{4\ell+5}\,$, (2.10)
where $C_{n\ell m}(a_{*})$ is a dimensionless, spin-dependent coefficient. For the dominant $\{2,1,1\}$ state [14]
$C_{211}(a_{*})\;\approx\;\frac{a_{*}}{48}\left(1-\frac{4\alpha^{2}}{a_{*}^{2}m^{2}}\right)^{\!1/2},$ (2.11)
which vanishes at the superradiance threshold $\alpha\to a_{*}m/2$ and is maximized at intermediate spin. The $\alpha^{4\ell+5}$ dependence makes higher-$\ell$ modes exponentially slower: at $\alpha=0.2$, the $\ell=2$ state grows $\sim\alpha^{4}\simeq 1.6\times 10^{-3}$ times more slowly than the $\ell=1$ dominant mode. Accordingly, the $\{2,1,1\}$ state is almost always the first to saturate, and it determines the timescale on which angular momentum is extracted from the BH.
@@
$\tau_{\rm SR}\;\equiv\;\Gamma_{n\ell m}^{-1}\;\approx\;5\,\mathrm{s}\left(\frac{M_{\rm BH}}{10^{-5}\,M_{\odot}}\right)\left(\frac{0.2}{\alpha}\right)^{\!9}\left(\frac{0.1}{C_{211}}\right),$ (2.12)
using $r_{g}/c\simeq 5\times 10^{-6}\,\mathrm{s}\,(M_{\rm BH}/M_{\odot})$ for the light-crossing time. Figure 2 shows $\tau_{\rm SR}$ in years as a function of $\alpha$ for modes $\ell=m=1,\ldots,4$; Fig. 2 shows the same rate in geometrized units $r_{g}\,\Gamma_{n\ell m}$, making the $\ell$-dependent peak structure visible.
@@
    Cutoff at the superradiance threshold. Both figures show $\Gamma_{n\ell m}\to 0$ as $\alpha\to m/2$, because the superradiant enhancement $C_{n\ell m}$ vanishes there. The instability therefore operates within a finite window $\alpha\in(0.05,m/2)$, peaking at $\alpha\sim 0.2$$0.3$.
@@ 3.2.1 Annihilation frequency and rate
$\Gamma_{a}\;\simeq\;C_{n\ell m}\left(\frac{\alpha/\ell}{0.5}\right)^{\!4\ell+4}\frac{G_{N}}{r_{g}^{3}}\,$, (3.13)
where the exponent $p\equiv 4\ell+4$ is determined by the power of $\alpha$ in the hydrogenic wavefunction overlap, and $C_{n\ell m}\approx 10^{-10}$ is a dimensionless coefficient that depends weakly on $n$ for $n\lesssim\ell+2$ [4, 9]. Concretely: for the lowest superradiant level $\ell=m=1$, one has $p=8$; for $\ell=m=2$, $p=12$; and so on. Since $\Gamma_{a}\propto\alpha^{4\ell+4}/r_{g}^{3}$ while $\Gamma_{\rm SR}\propto\alpha^{4\ell+5}/r_{g}$ and $\Gamma_{t}\propto\alpha^{9}/r_{g}^{3}$, the hierarchy $\Gamma_{\rm SR}\gg\Gamma_{t}\gg\Gamma_{a}$ holds over the entire parameter space of interest, confirming that annihilation is the slowest of the three processes.
@@ 6.2 Comparison to Related Work
The GW signatures of superradiant clouds around stellar-mass BHs are well studied [5, 4, 9, 19], with constraints on ultralight bosons from LIGO/Virgo spin measurements and non-detections of continuous GWs establishing the most robust exclusion regions in the $(\mu,M_{\rm BH})$ plane for $\mu\sim 10^{-13}$$10^{-11}\,\text{eV}$ [22]. The present work extends this program into qualitatively new territory in two respects.
@@
Second, the binary-perturbation analysis builds directly on the Baumann–Chia–Porto (BCP) Landau–Zener framework [7, 6] and the recent extension by Kyriazis & Yang [20], which derived the analytic GW strain template for tidally driven transitions. Previous applications of the BCP formalism focused primarily on the depletion of clouds (floating orbits, inspiral modifications, dephasing of the binary GW signal) in the context of LISA-band sources [7, 6, Zhang:2020], and on the resonance history in Ref. [29]. We are the first to apply the Kyriazis & Yang waveform template to PBH binaries in the MHz–GHz regime and to compare it directly to the sensitivity of an existing high-frequency detector, yielding a quantitative (negative) detectability result rather than a parametric estimate.
@@ 6.3 Caveats and Limitations
We have treated the isolated and binary channels independently, but in a realistic PBH population both may be active simultaneously or sequentially. Binary companions that sweep through resonance before the isolated cloud reaches saturation will partially deplete the cloud, suppressing the subsequent annihilation and level-transition signals [29]. Conversely, an isolated cloud that has already undergone significant annihilation before the binary resonance is crossed will produce a weaker binary-induced burst than our estimates suggest. A self-consistent treatment that evolves both channels simultaneously remains to be carried out.

### 10 Apr 2026 v3

@@ 1 Introduction
where $\Omega_{H}$ is the angular velocity of the Kerr horizon and $m$ the azimuthal quantum number, is exponentially amplified at the expense of the black hole’s rotational energy [35, 36, 26]. When the boson’s Compton wavelength is comparable to the gravitational radius, the amplified modes are gravitationally trapped and grow into a macroscopic quasi-bound cloud—a gravitational atom (GA) [13, 10]—with occupation numbers reaching $N\sim 10^{70}$$10^{80}$ [4]. The dynamics are controlled by the gravitational fine-structure constant
@@
The GW frequency is set almost entirely by $\mu$: annihilation of two cloud bosons into a graviton emits at $\omega_{\rm ann}\simeq 2\mu$, while level transitions between states with principal quantum numbers $n_{g}$ and $n_{e}$ emit at the Bohr frequency $\omega_{\rm tr}\propto\mu\alpha^{2}(n_{g}^{-2}-n_{e}^{-2})$. For stellar-mass BHs this places signals in the audio band probed by LIGO/Virgo/KAGRA, and an extensive literature has explored these GW signatures and the spin constraints they imply [10, 4, 20, 23]. A qualitatively distinct regime arises for primordial black holes (PBHs) [11, 27, 3]. For sub-solar mass PBH masses the superradiance condition selects boson masses $\mu\sim 10^{-7}$$10^{-5}$ eV, driving GW emission into the MHz–GHz band. This is precisely the frequency range targeted by resonant-cavity haloscopes such as ADMX, which were originally designed to search for axion dark matter but are equally sensitive to oscillatory strain at GHz frequencies, provided the signal persists for longer than the cavity ring-up time $\tau_{\rm ring}\sim 1/\Delta f_{\rm band}$. Also, notably, bosons in that mass range are independently motivated by the axion solution to the strong CP problem [24, 31, 32] and the cosmology thereof [25, 1, 15, 14].
@@
These signals have been studied in detail for isolated stellar-mass GAs [10, 4, 20], and we extend that framework here to the PBH mass range and the associated GHz frequency band. Bosenova collapses, in which the cloud implodes once its occupation number exceeds a critical value set by bosonic self-interactions, introduce additional burst-like GW emission and periodically reset the superradiant growth [33, 34]; we account for the resulting upper bound on cloud populations when deriving strain amplitudes.
In realistic astrophysical environments, however, GAs are often not isolated. A binary companion introduces a time-varying tidal potential that drives resonant transitions between GA levels whenever the orbital frequency sweeps through a Bohr frequency of the gravitational atom. In this regime the cloud behaves as a driven two-level system, and its response is captured by a Landau–Zener analysis [6]: depending on the adiabaticity parameter $z\equiv\eta^{2}/(|\Delta m|\dot{\Omega})$, where $\eta$ is the tidal mixing amplitude, transitions range from partial population transfer ($z\ll 1$) to complete level inversion ($z\gg 1$). The resulting GW signal is a short-duration transient burst whose characteristic strain and spectral shape have recently been computed by Kyriazis & Yang [21], whose formalism we adopt and apply to the PBH binary case. Binary-induced resonant transitions also imply:
@@ 2.2 Superradiant Instability Rates and Growth Timescales
$\Gamma_{n\ell m}=\frac{2r_{+}}{M}C_{n\ell}g_{\ell m}(m\Omega_{H}-\omega_{n\ell m})\alpha^{4\ell+5},\\$ (2.10)
with
$\quad C_{n\ell}=\frac{2^{4\ell+1}(n+\ell)!}{n^{2\ell+4}(n-\ell-1)!}\left(\frac{\ell!}{(2\ell)!(2\ell+1)!}\right)^{2},\\ g_{\ell m}=\prod_{k=1}^{\ell}\left[k^{2}(1-{a^{*}}^{2})+\left(a^{*}m-2r_{+}\omega_{n\ell m}\right)^{2}\right]$
[…]
$r_{+}=\frac{GM}{c^{2}}\left(1+\sqrt{1-\frac{a^{2}}{M^{2}c^{2}}}\right).$
@@
$\tau_{\rm SR}\;\equiv\;\Gamma_{n\ell m}^{-1},$ (2.11)
Figure 2 shows $\tau_{\rm SR}$ in years as a function of $\alpha$ for modes $\ell=m=1,\ldots,4$; Fig. 2 shows the same rate in geometrized units $r_{g}\,\Gamma_{n\ell m}$, making the $\ell$-dependent peak structure visible.
@@
    Cutoff at the superradiance threshold. Both figures show $\Gamma_{n\ell m}\to 0$ as $\alpha\to m/2$, because the superradiant enhancement $C_{n\ell}g_{\ell m}$ vanishes there. The instability therefore operates within a finite window $\alpha\in(0.05,m/2)$, peaking at $\alpha\sim 0.2$$0.3$.
@@ 3.2.1 Annihilation frequency and rate
$\Gamma_{a}\;\simeq\;K_{n\ell m}\left(\frac{\alpha/\ell}{0.5}\right)^{p}\frac{G_{N}}{r_{g}^{3}}\,$, (3.13)
where the exponent $p\equiv 4\ell+11$ is determined by the power of $\alpha$ in the hydrogenic wavefunction overlap, and $K_{n\ell m}\approx 10^{-10}$ is a dimensionless coefficient that depends weakly on $n$ for $n\lesssim\ell+2$ [4, 10]. Concretely: for the lowest superradiant level $\ell=m=1$, one has $p=8$; for $\ell=m=2$, $p=12$; and so on. Since $\Gamma_{a}\propto\alpha^{4\ell+11}/r_{g}^{3}$ while $\Gamma_{\rm SR}\propto\alpha^{4\ell+5}/r_{g}$ and $\Gamma_{t}\propto\alpha^{9}/r_{g}^{3}$, the hierarchy $\Gamma_{\rm SR}\gg\Gamma_{t}\gg\Gamma_{a}$ holds over the entire parameter space of interest, confirming that annihilation is the slowest of the three processes.
@@ 6.2 Comparison to Related Work
The GW signatures of superradiant clouds around stellar-mass BHs are well studied [5, 4, 10, 20], with constraints on ultralight bosons from LIGO/Virgo spin measurements and non-detections of continuous GWs establishing the most robust exclusion regions in the $(\mu,M_{\rm BH})$ plane for $\mu\sim 10^{-13}$$10^{-12}\,\text{eV}$ [23]. The present work extends this program into qualitatively new territory in two respects.
@@
Second, the binary-perturbation analysis builds directly on the Landau–Zener framework within [6] and the recent extension by Kyriazis & Yang [21], which derived the analytic GW strain template for tidally driven transitions. Previous applications of the Landau–Zener formalism focused primarily on floating orbits, inspiral modifications, dephasing of the binary GW signal, and the depletion of clouds in the context of LISA-band sources [6], and on the resonance history in Ref. [30]. We are the first to apply the Kyriazis & Yang waveform template to PBH binaries in the MHz–GHz regime and to compare it directly to the sensitivity of an existing high-frequency detector, yielding a quantitative (negative) detectability result rather than a parametric estimate.
@@ 6.3 Caveats and Limitations
We have treated the isolated and binary channels independently, but in a realistic PBH population both may be active simultaneously or sequentially. Binary companions that sweep through resonance before the isolated cloud reaches saturation will partially deplete the cloud, suppressing the subsequent annihilation and level-transition signals. Conversely, an isolated cloud that has already undergone significant annihilation before the binary resonance is crossed will produce a weaker binary-induced burst than our estimates suggest. A self-consistent treatment that evolves both channels simultaneously remains to be carried out.